find all the real zeros of F(x)=3x^4+4x^3-3x-4
Answer by stanbon(75887)

find all the real zeros of F(x)=3x^4+4x^3-3x-4
-----
Since the coefficients add up to zero, x = 1 is a root.
Use sythetic division:
1)....3....4....-3....-4
.......3....7.....4...|..0
----
Quotient: 3x^2 + 7x + 4
Solve to find other zeros:
3x^2 + 7x + 4 = 0
3x^2 + 3x + 4x + 4 = 0
3x(x+1) + 4(x+1) = 0
(x+1)(3x+4) = 0
---
x = -1 or x = -4/3
----
All the zeros: 1,-1,-4/3
==============================
Cheers,
Stan H.
===================